Write an equation of a line given the following information.
passes through the points (-2,6) and (-3,-4) can you explain each step please

Answer by Alan3354(21605)

Write an equation of a line given the following information.
passes through the points (-2,6) and (-3,-4) can you explain each step please
------------------------
The hard way:
1st, find the slope, m, of the line.
m = (diff in y)/(diff in x)
m = (-4 -6)/(-3 -(-2))
m = -10/-1
m = 10
------
Then, use either point in
y-y1 = m*(x-x1) where (x1,y1) is the point.
y-6 = 10*(x+2)
y-6 = 10x + 20
y = 10x + 26
That's it.
------------
The easy way:
|x y 1|
|-2 6 1| = 0
|-3 -4 1|
----------------
x*(6+4) - y*(-2+3) + (8 + 18) = 0
10x -y +26 = 0
y = 10x + 26